OpenStudy (anonymous):

[9.01] Which equation is a quadratic equation?

OpenStudy (goformit100):

quadratic equation (plural quad·rat·ic e·qua·tions) noun equation containing squared terms: an equation containing one or more terms raised to the power of two but no higher

OpenStudy (callisto):

@Karakuhnn I think the question is incomplete...

OpenStudy (anonymous):

a. 5x − 2 = y b. y + 15x^2 = (3x − 2)(5x + 7) C. y = 3x(5 − x) + 7 d.y + 2 = (3x^2 + 10)(x − 1)

OpenStudy (goformit100):

@JoannaiA

OpenStudy (anonymous):

A Quadratic Equation is one that has a degree of 2. Which of the choices has a squared term like \(x^{2}\)?

OpenStudy (anonymous):

B and D do

OpenStudy (anonymous):

Well, D indeed does have a squared term, but when you fully multiply out (3x^2 + 10)(x − 1), you get a cubic because (3x^2 + 10)(x − 1) = 3x^3 - 3x^2 + 10x - 10. See how it turns to a degree of 3? THerefore, it's B since it will stay with a degree of 2.

OpenStudy (anonymous):

oh ok i see..thank you!